# Assurance of AI Systems From a Dependability Perspective

### v2

@@ 4 Assurance for General-Purpose AI
Jacovi and colleagues [83] discuss “intrinsic” and “extrinsic” trust in AI systems. The latter is based on evaluation of observed behavior and we have previously, in Section 1, discussed the infeasibility of deriving high levels of assurance from observation alone. In those discussions, the obstacle was the vast number of tests needed for useful assurance in the absence of justified prior confidence. In the context of general-purpose AI systems, less assurance may be acceptable and, given some prior confidence, the number of tests might be feasible; instead, the difficulty becomes construction and evaluation of adequately wide-ranging tests, given that potential applications are unknown. Some developers of general-purpose tools are reported to manage this using crowdsourcing. [^29] For example, using the Amazon Mechanical Turk: https://www.mturk.com/. An alternative approach uses one AI system to test and evaluate another [66]. A reputation system, as sketched above, can be seen as a way of assigning and updating extrinsic trust in a “live” system.
In contrast to extrinsic trust, intrinsic trust is based on the ability of an AI system to explain or justify its behavior. There is a vast amount of work on “Explainable AI” and we mentioned counterfactual explanations and the fact that they can be manipulated in Section 3. Constructive explanations, an alternative to counterfactuals, provide reasons to persuade a “checker” that the AI system’s proposed action or output is appropriate, or at least reasonable. If the checker is a human, this requires them to be sufficiently expert in the application area that they can use the explanation to construct a chain of reasoning and verify its accuracy. This is unrealistic in many applications: for example, a lay person is unlikely to be able to verify an explanation of medical advice from an “AI doctor.” Recall also the problem discussed in Section 3 of generating and interpreting explanations when the world models of the AI system and the human checker are misaligned.
@@
The developers of LLMs and other general-purpose AI and ML systems are actively working to improve their trustworthiness [17]. We expect that we will see major and perhaps surprising developments in both capability and trustworthiness over the next few years, but we doubt this will achieve true assurance (i.e., explicit reasons for trust) and we believe that the methods presented here cover the feasible approaches for some degree of dependability assurance in applications of general-purpose systems. Because assurance for their general behavior is difficult, we recommend that developers of LLMs and other general-purpose systems provide APIs and “hooks” that those who construct applications on their foundation can use to program protections or guards suitable to their specific context.
@@ 5 Assurance and Alignment for AGI
Concerns raised by these developments, which are considered fanciful by some but inevitable by others, are that AGI poses “existential” threats to humanity ranging from widespread unemployment, destruction of social institutions, provocation of civil unrest or international conflict, all the way to our enslavement or extinction.
These concerns have become a topic of public discourse due to very rapid recent advances in the capabilities of systems using AI and ML. In the last ten years, AI systems based on Reinforcement Learning (RL) have beaten world champions in Chess and Go, outperformed conventional weather-prediction models, and solved the folding problem for proteins. These capabilities are rightly seen as impressive and beneficial but they excited little public debate. That step was achieved by OpenAI’s release of LLM-based ChatGPT in November 2022 [128], its adoption by Microsoft, and the release of similar systems by Anthropic, Google and others. [^30] See reports of the AI Index https://aiindex.stanford.edu. LLMs generate near human-quality text and code in response to modest “prompts” [127] and similar systems such as Stable Diffusion and Sora generate images and videos [114].
@@
In previous sections we have considered assurance for specific systems and for general-purpose tools based on AI and ML where the concern is that faulty behavior may lead to harm. For projected developments of current AI and ML systems and the potential emergence of AGI, however, the concern is not just faulty behavior but the social impact of new capabilities. In particular, an AGI system capable of setting its own goals might pursue objectives contrary to human interests. Notice that although the terminology is seldom employed in discussion of these topics, these are nonetheless dependability failures: an AI system with potentially contrary goals is a hazard that should be recognized and should be furnished with requirements to mitigate the danger, together with dependability requirements analysis to show that they do so. However, current practice in the field frames the assurance problem for AGI as ensuring that its goals align with those of human society [177]—and this needs to be maintained even though AGI may fall into the hands of bad actors, criminals, and adversaries.
@@
Some consider that “safe superintelligence is the most important technical problem of our time” (https://ssi.inc), but an insidious form of disruption may arise long before AGI is available: namely, AI systems that are good enough and cheap enough to displace (possibly superior) human services—what we might call Artificial Fairly Good Intelligence (AFGI). For example, well-researched journalism already finds it hard to compete with LLM-generated press releases and propaganda that masquerade as “news.” Related to this is another near-term disruption: the possibility that ubiquitous use of LLMs and related AI tools will reduce our collective sagacity by repeatedly circulating misinformation, falsehoods, and mediocrity, so that we lose the ability to access or recognize truth, novelty, or insight [116]. This extends to images, sound, and videos, where “deepfakes” enabled by diffusion models invert the adage that “seeing (and hearing) is believing.”
@@
Furthermore, LLMs and diffusion models are trained by scraping text, images, and all manner of other material from the Web (likely violating copyright in many instances [^31] New York Times, 27 December 2023.) and as they become widely used, their own output becomes a large part of the web corpus that trains the next generation and so on, potentially causing the Web to become full of LLM-generated pablum. There already is evidence for this: replying to a prompt asking it to tinker with malware, an LLM called Grok, developed by xAI, refused “as it goes against OpenAI’s use case policy.” OpenAI has nothing to do with xAI, so Grok presumably generated this response by scraping web text generated by OpenAI’s LLMs. This phenomenon is termed Model Collapse [157] and is predicted to “dumb down” the overall information content of the Web. Meanwhile, recommender systems for music, video, and books may lead us to serendipitous and enjoyable discoveries, but they can also be manipulated to channel our attention along predetermined paths.
@@
A generally anticipated disruption is widespread unemployment when capabilities a little better than those available today, but still far short of AGI, displace office jobs ranging from clerks to middle-management, software coding, routine design, and professions such as lawyer and doctor. These AFGI capabilities might be superior to human performance, or inferior but cheaper and still “good enough”; either way, they could displace human jobs. Beyond harm to individuals, unfettered use of AFGI, particularly in conjunction with polarization perpetrated by social media, could undermine our institutions, or our trust in these. All these disruptions could be chronic rather than acute: that is, they could develop over a period (e.g., creeping unemployment) rather than abruptly and, as with climate change, this may make recognition difficult, and effective response hard to mobilize.
@@
In addition to these cases, where harm may be unintended, there is the threat that AFGI could multiply the capabilities of adversaries and hackers (e.g., by automating the coding of viruses, or the generation of phishing emails and provocative social media posts). In the hands of adversaries, AFGI-enabled disinformation campaigns and social media manipulations could sow societal discord and conflict. Until recently, interventions of this kind required the capabilities and resources of a nation state, but are now coming within reach of lone actors. Then there are truly dystopian applications of AFGI such as “Lethal Autonomous Weapons Systems” (LAWS, aka. “killer robots”). Most countries have policies that require some form of human control or approval over these (e.g., USA [156]) but it is unknown how these limitations will hold up in combat. And there are AI systems that contribute to lethality without being weapons. For example, AI-powered intelligence systems identify vastly more potential targets more quickly than humans can (e.g., 100 a day vs. 50 a year [9]). Humans are required to select and approve actual targets from those identified but the sheer numbers make meaningful oversight difficult.
@@
A very short“consensus” paper with 24 distinguished authors outlines several other potential risks of AI and proposes various technical and governance measures [16]. In that regard, the United States Government recently (30 October 2023) issued Executive Order 14110 on the “Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence.” The Executive Order directs the National Institute of Standards and Technology (NIST) to “develop guidelines and best practices to promote consensus industry standards that help ensure the development and deployment of safe, secure, and trustworthy AI systems” (see https://www.nist.gov/aisi). Similarly, the European Union has a recent (May 2024) law on AI that takes a risk-based approach to AI applications but also, for the first time, proposes to regulate the underlying technology of foundation models and General Purpose AI (GPAI). Foundation models must comply with transparency obligations, and “high-impact models with systemic risk” will have to conduct model evaluations, assess and mitigate systemic risks, conduct adversarial testing, report to the European Commission on serious incidents, ensure cybersecurity and report on their energy efficiency. GPAIs with systemic risk may rely on codes of practice to comply with the new regulation.
@@
These steps are well-intentioned but apart from the EU’s GPAI proposals, we consider them unlikely to provide much benefit. Every societal hazard posed by recent computer systems was unanticipated and unrecognized until the harm was done, and subsequent regulation has proved ineffective or counterproductive. For example, the advent of advertising as the main revenue source for Web services has led to universal surveillance of personal online activity, to the extent that even our cars now invade our privacy. [^32] See https://foundation.mozilla.org/en/privacynotincluded/categories/cars/. This was not anticipated and regulatory responses such as “cookie warnings” are a source of annoyance rather than protection, and indicate technological illiteracy on the part of regulatory agencies and lawmakers. Similarly, the poisonous impact of “algorithmic” traffic generation on social media was not anticipated and still has no effective response.

### v3

@@ 4.4 Reputation
4.5 Checkable Outputs
Jacovi and colleagues [116] discuss “intrinsic” and “extrinsic” trust in AI systems. The latter is based on evaluation of observed behavior and we have previously, in Section 1.1, discussed the infeasibility of deriving high levels of assurance from observation alone. In those discussions, the obstacle was the vast number of tests needed for useful assurance in the absence of justified prior confidence and the analysis required to justify extrapolation from past to future behavior. In the context of general-purpose AI systems, less assurance may be acceptable and, given some prior confidence, the number of tests might be feasible; instead, the difficulty becomes construction and evaluation of adequately wide-ranging tests, given that potential applications are unknown. Some developers of general-purpose tools are reported to manage this using crowdsourcing. [^29] For example, using the Amazon Mechanical Turk: https://www.mturk.com/. An alternative approach uses one AI system to test and evaluate another [92]. A reputation system, as sketched above, can be seen as a way of assigning and updating extrinsic trust in a “live” system.
In contrast to extrinsic trust, intrinsic trust is based on the ability of an AI system to explain or justify its behavior. There is a vast amount of work on “Explainable AI” and we mentioned counterfactual explanations and the fact that they can be manipulated in Section 3. Constructive explanations, an alternative to counterfactuals, provide reasons to persuade a “checker” that the AI system’s proposed action or output is appropriate, or at least reasonable. If the checker is a human, this requires them to be sufficiently expert in the application area that they can use the explanation to construct a chain of reasoning and verify its accuracy. This is unrealistic in many applications: for example, a lay person is unlikely to be able to verify an explanation for medical advice from an “AI doctor.” Recall also the problem discussed in Section 3 of generating and interpreting explanations when the world models of the AI system and the human checker are misaligned.
@@ 4.5 Checkable Outputs
4.6 Guardrails and Architecture Patterns
The developers of LLMs and other general-purpose AI and ML systems are actively working to improve their trustworthiness [21, 23]. However, as we will describe in the following section, much of the response has concerned “existential” risks rather than everyday dependability (or misalignment vs. reliability [24]). This is unfortunate as we think that dependability is the more urgent problem and is becoming even more so as AI applications built on general purpose foundations become widespread due to rapid improvements in the performance of AI foundation models. Ironically, we fear these improvements may lead to less dependable applications as those who build applications around foundation models gain trust in them and reduce their external checks and supervisory systems, instead building them as “guardrails” on the foundation model itself.
Referring back to the architecture diagrams of Section 2.5, the minimal arrangement there labeled 1 can be recast as follows when an LLM or other foundation model is employed.
The idea is that the perception, local model, and action elements are all now part of the LLM. Thus, when we add “guardrails” similar to the earlier Architecture 3, they are likewise internal as portrayed below.
Like that earlier architecture 3, the guardrails are driven from the same perception and local model as the main action system and provide no protection for faults in those elements. But unlike the earlier architecture, here it is difficult to argue that the internal guardrails are either independent or diverse. Furthermore, there is evidence that foundation models are able to fake their alignment [91, 176]. This is not to assert that internal guardrails cannot provide benefit, but that the benefit cannot be assured.
We expect to see further major and perhaps surprising developments in both capability and superficial trustworthiness of LLMs and other general-purpose foundation models over the next few years, but we doubt these will deliver true assurance: minor failures may become less frequent as the foundation models improve, but there is little reason to suppose that serious ones will decline significantly. [^30] This is due to “fat tails” on the failure distribution. Koopman provides an example: in a variant with thin tails, we might have 100 serious faults each with an arrival rate of 1 in 100 million demands; in a thick-tailed variant we might have 100,000 serious faults each with an arrival rate of 1 in 100 billion demands. Both variants expect a bad event every million demands but after exposure or assurance equivalent to a billion demands we will have seen all the faults of the first variant but only 1% of the second. Instead, we must continue to look to external checks as portrayed below, which is based on Architecture 5 of Section 2.5.
We believe that the methods presented here cover the feasible approaches for some degree of dependability assurance in applications of general-purpose systems. In order to invoke external checks for runtime assurance, we recommend that developers of LLMs and other general-purpose systems provide APIs and “hooks” that those who construct applications on their foundation can use to program protections and to invoke external guards suitable for their specific context. These could include human review, automated review by a diverse model, checking against normative rules, and adjustment in response to reputation scores. We note that some AI frameworks already provide APIs for accessing external reasoning tools (e.g., [165]) and it is possible that these and also mechanisms for chain of thought reasoning could be used or adapted for the assurance purposes advocated here.
We conclude this section by revisiting the traditional strategies for dependability.
Fault avoidance.
Although there is much merit in making the general-purpose AI as reliable and trustworthy as possible, it is impossible to provide assurance without knowing the application context and potential system hazards.
Fault tolerance.
Generic guardrails can be seen as elements in a fault avoidance strategy. But guardrails and other checks constructed later in the development process, when the system context is known, can be part of a fault-tolerance strategy for systems built around general-purpose AI. For credible assurance, these components should be external to the main AI so that claims of independence and diversity are plausible (recall the third architecture diagram above).
Failure Management and Resilience.
The issues and challenges discussed in this section illustrate the value of resilience strategies. General-purpose AI may become widely deployed in the absence of context-specific protections, simply on the (inappropriate) basis of interpersonal trust. Resilience can provide some measure of risk management and social trust. We note that a wide range of potential harms must be considered and anticipated, including those of existing and possibly regulated risks as well as general societal harms.
@@ 5 Assurance and Alignment for AGI
Concerns raised by these developments, which are considered fanciful by some but inevitable by others, are that AGI poses “existential” threats to humanity ranging from widespread unemployment, destruction of social institutions, provocation of civil unrest or international conflict, all the way to our enslavement or extinction. Separate from threats posed by AGI itself are concerns that near-AGI (so-called “frontier” models and systems) could enable bad actors to develop catastrophic cyber, chemical, biological, radiological, and even nuclear (CBRN) threats and attacks. [^31] The necessary information is assumed to be present on the Web, but dispersed and not in a form that is useful to outsiders; but frontier LLMs may be able to systematize, summarize, and synthesize this information into a readily exploitable form.
These concerns have become a topic of public and government discourse due to very rapid recent advances in the capabilities of systems using AI and ML. In the last ten years, AI systems based on Reinforcement Learning (RL) have beaten world champions in Chess and Go, outperformed conventional weather-prediction models, and solved the folding problem for proteins. These capabilities are rightly seen by scientists as impressive and beneficial (e.g., 2024 Nobel Prizes for physics and chemistry) but they excited little public attention. That step was achieved by OpenAI’s release of LLM-based ChatGPT in November 2022 [184], its adoption by Microsoft, and the release of similar systems by Anthropic, Google and others. [^32] See reports of the AI Index https://aiindex.stanford.edu. LLMs generate near human-quality text and program code in response to modest “prompts” [182], and similar systems such as Stable Diffusion and Sora generate images and videos [164].
@@
In previous sections we have considered assurance for specific systems and for general-purpose tools based on AI and ML where the concern is that faulty behavior may lead to harm. For projected developments of current AI and ML systems and the potential emergence of AGI, however, the concern is not just faulty behavior, nor even system failures, but the social impact of new capabilities. In particular, an AGI system capable of setting its own goals might pursue objectives contrary to human interests. Notice that although the terminology is seldom employed in discussion of these topics, these are nonetheless dependability failures and can be examined from that perspective: an AI system with potentially contrary goals is a hazard that should be recognized and should be furnished with requirements to mitigate the danger, together with dependability requirements analysis to show that they do so, and a system boundary and implementation that assuredly applies them. However, current practice in the field frames the assurance problem for AGI as ensuring that its behavior and goals align with those of human society [244, 84, 119, 24]—and this alignment needs to be maintained even though AGI may fall into the hands of bad actors, criminals, and adversaries.
5.1 Fairly General/Good AI, AFGI
@@
Some consider that “safe superintelligence is the most important technical problem of our time” (https://ssi.inc), but an insidious form of disruption may arise long before AGI is available: namely, AI systems that are good enough and cheap enough to displace (possibly superior) human services—what we might call Artificial Fairly General (or Fairly Good) Intelligence (AFGI, pronounced “AFF-GEE”). For example, well-researched journalism already finds it hard to compete with LLM-generated press releases and propaganda that masquerade as “news.” Related to this is another near-term disruption: the possibility that ubiquitous use of LLMs and related AI tools will reduce our collective sagacity by repeatedly circulating misinformation, falsehoods, and mediocrity, so that we lose the ability to access or recognize truth, novelty, or insight [168, 183]. This extends to images, sound, and videos, where “deepfakes” enabled by diffusion models invert the adage that “seeing (and hearing) is believing.” Beyond the hazards of these somewhat passive, human-directed applications of AI and ML are those of active “agents” driven by similar technology. These act autonomously, performing self-determined tasks to meet human-determined goals. There are many possibilities for failure here, including poorly selected tasks, misinterpreted goals, and a misperceived environment.
In the mere six months since the first edition of this report was issued, AFGI capabilities have developed at a remarkable rate. One of the more significant developments is the “chain of thought” (CoT) approach to “reasoning” where a problem is broken in smaller chunks that are solved sequentially [251]. This was first widely seen in OpenAI’s o1 model, but is now commonplace (giving rise to the terminology “Large Reasoning Model,” or LRM). Another is use of “distillation” where a powerful LLM is used to improve and fine tune a weaker model [102]. This can generate models with performance close to the powerful LLM but small enough to run on a single modest computer or even a mobile device. Somewhat related is use of fewer bits to quantize weights and parameters in the underlying neural nets: e.g., using $\{-1,0,+1\}$ (1.58 bits) instead of 4, 8, or 16 bit integers [166]. Again, the reduced model is much smaller but almost as good as the original and has less need of costly high-performance GPUs. Finally, the open source DeepSeek R1 model that was released in January 2025 achieves performance comparable to OpenAI o1 at a fraction of the development cost (so it is claimed). [^33] Contrary claims assert it is a distillation of other models. And in addition to LLMs, AI corporations and cloud providers now supply components and environments for constructing AI agents. [^34] For example, see https://aws.amazon.com/what-is/ai-agents/.
If we project these advances and general research progress forward a few years, we can expect very powerful and affordable AFGI systems to be widely available and we propose that it is important to consider the risks they pose in addition to those of speculated AGI systems. (It is a controversial topic whether AFGI systems lead to AGI: there is an empirical “scaling law” [126] that LLMs get better the bigger they are. Some believe this leads inevitably to AGI, others that (much) more is required and that LLMs are already approaching the limits of scaling. We subscribe to the latter opinion because AGI surely requires the ability to reason over models of the world and, as we have stressed several times, LLMs are “model free.”)
The interim report from the Seoul conference on the safety of advanced AI identifies a broad range of risks[21]. Unfortunately, the report explicitly chooses not to address “Narrow AI,” which is “used in a vast range of products and services…and can pose significant risks in many of them” and focuses on frontier LLMs. And even within that coverage, the “Seoul Commitments” focus on “existential risks” initiated by malign users with access to hypothesized models with near-AGI capabilities. [^35] The Seoul Report states that its restricted focus is due to “the limited timeframe for writing this interim report.” The recent full report [24] has much broader coverage. Following the Seoul Conference, many corporate developers of frontier AI systems signed up to the Seoul Commitments with their focus on existential risks and have published accounts of progress detecting and controlling these. [^36] A list of corporate safety policies is available at https://metr.org/faisc. In addition, several countries established oversight bodies (e.g., the UK AI Safety/Security Institute, aisi.gov.uk) and these also tend to focus on existential risks.
We, on the other hand, believe that the capabilities of current or near-term AFGI have reached a stage where they (or systems based on similar technology) may be deployed in “narrow” applications (as discussed in Sections 3 and 4) in preference to custom solutions, and hence the safety risks in these applications should be considered part of advanced AI safety. This means we must consider component failures initiated by faults within the system (e.g., “hallucinations”) and “normal accidents” due to system failures, as well as those initiated in the environment (e.g., by rogue users), and we should consider general dependability and creeping disability as well as existential risks.
We examined dependability in applications and in general-purpose models such as LLMs in the previous two sections and we do not think availability of AFGI frontier models changes our assessments, other than the scale and urgency of the risk—although we are concerned that the apparent quality of these models will encourage construction of “guardrails” on or within the frontier model itself, rather than externally, where diversity and independence can provide rather more—but still modest—assurance. And we are also concerned that the fluency of AFGI systems (for images, speech, and video, not just language) will embed them deeply into social contexts so that the system boundary is unknown and low-level system failures may become endemic and serious ones become possible.
@@
Between failures and existential risks there is a category of societal harms and disruptions that AFGI may precipitate. First, there are unfortunate consequences precipitated by the training of frontier models. LLMs and diffusion models are trained by scraping text, images, and all manner of other material from the Web. This likely violates copyright in many instances [^37] New York Times, 27 December 2023. and jeopardizes the livings of writers, artists, and musicians and their associated ecosystems as LLMs plagiarize their work. Furthermore, as LLMs become widely used, their own output becomes a large part of the web corpus that trains the next generation and so on, potentially causing the Web to become full of LLM-generated pablum. There already is evidence for this: replying to a prompt asking it to tinker with malware, an LLM called Grok, developed by xAI, refused “as it goes against OpenAI’s use case policy.” OpenAI has nothing to do with xAI, so Grok presumably generated this response by scraping web text generated by OpenAI’s LLMs. This phenomenon is termed Model Collapse [220] and is predicted to “dumb down” the overall information content of the Web. Meanwhile, search and recommender systems for music, video, and books may lead us to serendipitous and enjoyable discoveries, but they can also be manipulated to channel our attention along predetermined paths.
Another consequence of training is massive use of electric power and water resources by the servers employed, potentially harming the environment. Government and corporate initiatives are proposing to spend huge sums on development and use of additional server farms (the “Stargate” project targets half a trillion dollars,383838see https://openai.com/index/announcing-the-stargate-project/. or approximately the GDP of Norway or Malaysia). In addition to environmental harm, the costs involved may limit development, and ultimately control, of frontier technology to just a few countries and corporations, so that an oligarchy of magnates could subvert the structure of established societies and their governments.
@@
Moving from training to deployment, a generally anticipated disruption is widespread unemployment or other labor market disruption when AFGI capabilities little better than those available today displace office jobs ranging from clerks to middle-management, together with skilled jobs such as software coding and routine design, and professions such as lawyer and doctor. These AFGI capabilities might be superior to human performance, or they might be inferior but cheaper and “good enough”; either way, they could displace human jobs. Beyond harm to individuals, unfettered use of AFGI, particularly in conjunction with polarization perpetrated by social media, could undermine our institutions, or our trust in these. All these disruptions could be chronic rather than acute: that is, they could develop over a period (e.g., creeping unemployment) rather than abruptly and, as with climate change, this may make recognition difficult, and effective response hard to mobilize.
@@
In addition to cases where harm may be unintended, there is the threat that AFGI could multiply the capabilities of adversaries and hackers (e.g., by automating the search for vulnerabilities and the coding of viruses to exploit them, or the generation of phishing emails, provocative social media posts, and hurtful “deepfake” images and videos). In the hands of adversaries, AFGI-enabled disinformation campaigns and social media manipulations could sow societal discord and conflict. Until recently, interventions of this kind required the capabilities and resources of a nation state, but are now coming within reach of small groups or lone actors.
Then there are truly dystopian applications of AFGI such as “Lethal Autonomous Weapons Systems” (LAWS, aka. “killer robots”). Most countries have policies that require some form of human control or approval over these (e.g., USA [217]) but it is unknown how these limitations will hold up in combat. And there are AI systems that contribute to lethality without being weapons. For example, AI-powered intelligence systems identify vastly more potential targets more quickly than humans can (e.g., 100 a day vs. 50 a year [11]). By doctrine, humans are required to select and approve actual targets from those identified, but the sheer numbers make meaningful oversight difficult.
A very short “consensus” paper with 24 distinguished authors outlines several other potential risks of AI and proposes various technical and governance measures [22]. In that regard, the UK held the first AI Safety Summit at Bletchley Park in 2023 and established an AI Safety Institute (AISI) that performs “rigorous AI research to enable advanced AI governance” (https://aisi.gov.uk). Likewise, on 30 October 2023 the United States Government issued Executive Order 14110 on the “Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence.” The Executive Order directed the National Institute of Standards and Technology (NIST) to “develop guidelines and best practices to promote consensus industry standards that help ensure the development and deployment of safe, secure, and trustworthy AI systems” (see https://www.nist.gov/aisi), but this order was revoked by the successor government on 20 January 2025. The UK subsequently renamed its AISI as the AI Security Institute, and the UK and the USA declined to endorse a declaration on “inclusive and sustainable” AI that was signed by 58 other countries at the February 2025 AI Action Summit in Paris.
@@
The European Union has a recent (May 2024) law on AI that takes a risk-based approach to AI applications, categorizing them in four levels: minimal, limited, high, and unacceptable. In addition, for the first time, it proposes to regulate the underlying technology of foundation models and General-Purpose AI (GPAI). Foundation models must comply with transparency obligations, and “high-impact models with systemic risk” will have to conduct model evaluations, assess and mitigate systemic risks, conduct adversarial testing, report to the European Commission on serious incidents, ensure cybersecurity and report on their energy efficiency. GPAIs with systemic risk may rely on codes of practice [^39] See https://artificialintelligenceact.eu/introduction-to-codes-of-practice/. to comply with the new regulation.
@@
Most of these steps seem well-intentioned [^40] A contrary point of view is that the focus on existential risks is a form of regulatory capture (by distraction), allowing AI corporations unregulated freedom to perpetrate other risks [129]. but apart from the EU’s GPAI proposals, we consider them unlikely to provide much benefit. Every novel systemic societal hazard posed by recent computer systems was unanticipated and unrecognized until the harm was done (they can be seen as system failures), [^41] Merton [177] was the first to explicitly identify unanticipated consequences, which later became more often termed “unintended” consequences; Zwart [254] criticizes this transition. and subsequent regulation has proved ineffective or counterproductive. For example, the advent of advertising as the main revenue source for Web services has led to universal surveillance of personal online activity, to the extent that even our cars invade our privacy. [^42] See https://foundation.mozilla.org/en/privacynotincluded/categories/cars/. This was not anticipated and regulatory responses such as “cookie warnings” are a source of annoyance rather than protection, and indicate technological illiteracy on the part of regulatory agencies and lawmakers. Similarly, the poisonous impact of “algorithmic” traffic generation on social media was not anticipated and still has no effective response.
5.2 Resilience as a Key Response to AFGI/AGI
The emergence of AFGI and potentially of AGI requires an eco-system for impact monitoring including citizen groups, government agencies such as AISI, international NGOs, and risk owners and their regulators. We suggest that, in addition to monitoring AI technology, these entities should monitor its impact on society, so that unanticipated risks and incipient system failures can be detected early, and addressed. Monitored risks should include chronic and creeping developments such as accumulation [127] and gradual disempowerment [147]. In addition to technological adjustments and assurance, risks should be mitigated through societal adaptation for, as we have already noted several times, safety and assurance are system properties and with deployment of AFGI the system can extend widely into its societal context. Thus, in parallel with mechanisms to protect against the risks of AFGI, we also need measures to make our systems and society more resilient to those risks. So, for example, if the hazard of unemployment is to be mitigated, there should be research and action on how humans and AI can productively work together, with widespread training on best practices.
We also suggest that hazard analysis should be performed according to best practices (which will need further development to deal with the unique hazards of AI) and that the upper levels of risk should be more finely graduated than simply “existential.” True existential risk refers to the extinction of humanity or civilization as we know it, comparable to nuclear war. There are many levels of serious and catastrophic risk below this and they require graduated levels of monitoring and assurance, with special care to ensure they do not accumulate. (Although it is focused on climate change, a report of the IPCC provides useful general descriptions and definitions for extreme risk and risk management together with resilience strategies such as coping, incremental adjustment, transformation, and adaptation [81].) Catastrophic system failures often accumulate from a combination or succession of smaller faults and abnormalities within a complex system (recall Perrow’s “normal accidents”) and safety is best achieved by making systems more resilient through being able to break and work around failure pathways (e.g., by avoiding what Perrow calls “tight coupling” and “interactive complexity” [185]) rather than striving to eliminate all component and subsystem failures.
Furthermore, it does not seem widely appreciated in the AI community that assurance does not scale linearly with risk: the cost and effort and the techniques employed for assurance of critical avionics, for example, are totally different than those for less critical systems, to the extent that the assurance for a Level A avionics system typically costs more than development of the system. Yet we see assessment for CBRN hazards in LLMs performed by “red teams.” This may suitable of initial exploration, but assurance against the asserted magnitude of this risk requires vastly more effort (and/or societal adaptation, such as surveillance, controlled access to precursors, development of protective measures, antidotes, hardening of targets, and so on).